The minute hand of a wall clock is 14 cm long. In the time interval from 6 PM to 6 : 30 PM, the tip of the
minute hand has:
(A) average velocity = 0.56 m/hr
(B) average speed = 0.88 m/hr = 0.88 m/hr)
(C) average acceleration = 3.52 m/hr
D) travelled distance = 0.44 m

Grade:12th Pass

1 Answers

vikas askiitian expert
13 years ago

average velocity = total displacement /total time

total displacement = 2L = 28cm =0.28m

total time = 30mins = 30/60 hr = 0.5hr

average velocity = 0.28/0.5 = 0.56km/hr                                ansA

 

average speed = total distance/total time

total distance = piL = 0.4396m

total time = 30mins = 0.5hr

average speed = 0.4396/0.5 = 0.88m/hr                               ansB

 

distance travelled = piR = 3.14*14 = 43.99cm=.44m              ansD